### v8

# Aristotle’s Second-Order Logic

Abstract
This paper has two aims. The first aim is to argue that an extension of second-order logic in its natural deduction presentation is the most adequate framework to formalize the metalogic of the Analytics and the logic implicit in the Topics as well as the reasoning present in other works of Aristotle. The second aim is to provide evidence for the thesis (defended by Bobzien[4]) that ancient logic had sufficient means to deal effectively with multiple generality.
Introduction
In the last decades there has been an increased interest in the interpretation of Aristotle’s Organon using the framework of modern mathematical logic, the pioneers of this approach being found, for instance, among the Polish school of logicians and philosophers of the first half of the twentieth century. Similar research has also been carried out for other schools of ancient logic, notably for Stoic logic and ancient Indian logic. In this context some researchers have argued that modern and contemporary formal and philosophical approaches to logic were already present among ancient logicians. Thus we go beyond merely using modern formal structures and methods as interpretative tools but seek evidence (as far as the unfortunately fragmentary state of the material is concerned) that concepts and structures very similar to these were already present therein [^2] Some arguments for the legitimacy of deploying modern formal concepts to interpret historical philosophical systems (what is called the hermeneutical ’virtuous circle’) are found in [10]. We add here that the ability of an ancient philosophical text to be translated in a more-or-less direct way into a variant of a standard modern formal system is in itself remarkable. From the start the choice of the most adequate system is not obvious and the situation is usually one of great (even unsurmountable) difficulties.. For example in the Prior Analytics Aristotle seems to make the distinction between a term and its extension, between proof theoretic arguments and and model theoretic concepts such as soundness by making use of the construction of models to serve as counterexamples. The modal syllogistic can be axiomatized and given a possible world semantics[14]. Stoic logic can be interpreted as a sequent calculus for propositional logic - in fact as a substructural relevance logic[4].
The present work’s goal is to look at the metalogic (the reasoning in the arguments Aristotle gives, for example, about the conversion of propositions or the validity and non-validity of syllogisms) employed in the Analytics and the logic implicit in the Topics and to argue that: i) it corresponds to an extension of second-order logic - presented in Jaśkowski333in this paper when we refer to natural deduction we assume that the Jaśkowski style presentation is meant. Contrary to Gentzen’s sequent-style presentation this version resembles more closely how actual reasoning is carried out. We also deploy its linearized version (also called Lemmon style). natural deduction style - and many of Aristotle’s statements and proofs involve nested quantifiers and non-monadic predicates, ii) that Aristotle was conscious of the natural deduction rules and precise in their application, iii) to give a first sketch of a formal reconstruction and interpretation of the system of argumentation (in which definitions or other kinds of hypothesis were tested and possibly refuted) for which the Topics was originally written.
The paper is organized as follows. In section 2 we focus on the metalogic of the Analytics and show that its most natural interpretation is as an extension of second-order logic which we call Aristotelian second-order logic (ASOL). ASOL is second-order logic extended with a collection of second-order $n$-ary predicates having as arguments predicate terms and possibly individual terms. We also include some remarks on how the formalization of the modal syllostic in [14] can be embedded into ASOL. In section 3 we argue that Aristotle consciously employs natural deduction rules for universal and existential quantification. In the Analytics the key second-order binary predicates are posited to be definable (at least in the non-modal setting) in terms of strict second order logic, which expresses Aristotle’s proto-Fregean extensionalism. In section 4 we study some key aspects of Aristotle’s Topics and show that ASOL is the best choice for representing the underlying logic. Here we require a richer collection of second-order predicates (expressing the predicabilia) as compared to the Analytics; we do not have any ’extensionalist’ reduction of the main second-order predicates to pure second-order logic, more specifically, to relations between first-order predication. A topic is a sentence ASOL in general universally quantified over most of the monadic predicate variables occurring in the sentence. These universally quantified monadic predicate variables are then supposed to be instantiated according to the dialectical needs at hand. Although the structure of the Topics consists mainly in a rather tedious enumeration of topics, the most purely logical among them can be organized into groups wherein all the topics in a group are logical consequences of a certain paradigmatic topic. Often Aristotle indicates explicitly that for two closely related topics (mentioned in succession) one is the consequence of the other. This circumstance provides evidence to gauge the strength and nature of the logic implicitly present. In section 5 we look at book V which deals with property (idion) and discuss some of the challenges facing a formalization of the topics included therein as well as analyzing Aristotle’s treatment of superlatives in book VII, 152a.
In the concluding section we touch upon the significance of our approach for the history of the development of logic and in particular the connection between this interpretation of the Organon and Frege’s logic. We end with a discussion about how the system of debate for which the Topics was written might be reconstructed and interpreted according to ASOL.
In order to provide more context for section 4 in appendix 1 we list a formalization of an important part of the topics of book IV, focused on the relation between genus and species. In appendix 2 we illustrate the dialectical process for which the Topics was written by exhibiting a ASOL formalization of a key section of Plato’s Lysis. Appendix 3 presents a minimal set of axioms for a formalization of the first proposition of Euclid and appendix 4 includes the definition of ASOL and its system of natural deduction. This paper does not assume any technical knowledge of modern formal logic beyond an understanding of the language and rules of (extended) second-order logic presented in appendix 4.
Metalogic of the Analytics
In this paper we will work in Aristotelian Second-order logic (ASOL) which is standard second-order logic extended with a collection of second-order $n$-ary predicates having as arguments predicate terms and possibly individual terms. We consider its standard natural deduction presentation. See appendix 4 for details. In this paper we consider only classical rather than intuitionaistic ASOL, that is, we have the rule $\bot_{c}$.
There have been many attempts to formalize Aristotle’s syllogistic expounded in the Prior Analytics in both its strictly assertoric and modal variants (see [14] for references). Some approaches involve monadic (modal or otherwise) first-order logic and other approaches use variants (modal of otherwise) of propositional logic. Additionally many authors assume that Aristotle reasons much like a modern logician by constructing models to refute the validity of certain syllogistic figures. The above approaches usually set up a minimal set of axioms together with a deductive system and attempts to show the agreement (at in least for most cases) between the derivable syllogisms and those considered valid by Aristotle (how Aristotle himself derived or justifies the validity or non-validity playing a secondary role). Our concern is different. We are interested in the structure of the proofs in the Analytics themselves, how Aristotle argues for the validity or non-validity of a certain figure or propositional conversion. The modern division between metalogic and object logic does not seem to be, at first glance, so evident here. Aristotle does not simply posit some axioms and a deductive system and show how some figures are valid and some are not (with the use of models). Rather it appears that what we now call metalogic and object logic form an inextricable whole. So the best formalization must be directly about this whole, about the logical and argumentative structure of the Analytics themselves. A noteworthy characteristic of the Analytics is the presence of variable letters ranging over Aristotelian terms which (so we argue) usually correspond to monadic predicate terms in second-order logic. There is a distinguished class of entities for which it is meaningful to ask if such entities fall or do not fall under such monadic predicate term - these we call individuals. Thus we can ask of a monadic predicate term is predicated or not of a given individual.
It is patent that a formalization of Aristotle’s reasoning must be a term-logic, that is, terms must be the basic building blocks - the most basic division being between individual terms (corresponding to ’first substances’) and non-individual terms. Sentences (for the moment we consider here only ’atomic’ sentences (heis logos, De Int. 17a16-17) are then built up from combinations of terms in a specified manner. Predication - and in Aristotle there are many distinct notions of predication hiding under the verb esti - corresponds to combinations of terms tagged according to the kind of predication being considered. Thus if ① corresponded to predicating a non-individual term $A$ of an individual term $a$ (let us say accordingo to ’essence’) and ② corresponded to predicating a non-individual term $B$ of another $A$ according to ’essence’ then grammatically (for SOV languages like Ancient Greek) and notationally the most natural representation would be aA① and AB②. Such a representation is necessarily if we wish to formalize correctly’s Aristotle’s essentially intensional logic: the predication of definition and the predication of property (idion) both imply that the terms have equal ’extension’ (in the sense of individuals falling under them) but the types of predication are different as are the terms: ’being able to laugh’ is not the same term as ’rational animal’ but both are co-extensional with ’man’. Thus we have grounds for holding that an extension of second-order logic like ASOL (in which the circled number tags correspond to our second-order predicates and ① to simple concatenation in reverse order $Aa$) is the most adequate minimal formalization for the logical-argumentative apparatus in the Organon and even, as we shall discuss later, in the Physics. What we wrote above applies to ’atomic’ sentences. But do not modern formal systems with their full inductive definitions of formulas have greater complexity and precision than can be reconstructed from Aristotle’s system embodying an only seemingly rudimentary account of complex sentences ? Against this worry we reply that the fact of the matter is that rigorous structural inductive definitions of sentences are not an exclusive possession of modern logic. A perfectly ’modern’ definition of the formulas of what we would now consider an extension of modal propositional logic (a multi-modal propositional logic) is found in Ockham’s Summa Logicae[11][79-80]. Going back to antiquity there is strong evidence for such inductive definitions in Stoic Logic[4].
The four basic kinds of protasis dealt with in the Analytics can be seen as corresponding to four distinct second-order binary predicates, even if we grant that such second-order predicates can be defined without such second-order predicates. Syllogisms correspond very naturally to complex sentences in ASOL of a certain structure:
$\forall A,B,C(\phi(A,B)\&\psi(B,C)\rightarrow\rho(A,C))$
where $\phi,\psi$ and $\rho$ range among certain atomic second-order binary predicates (the order of the free variables is not necessarily the order of the free variables in the atomic binary predicate). For example, the Barbara syllogism corresponds to:
$\forall A,B,C(\kappa AB\&\kappa BC\rightarrow\kappa AC)$
where the second-order atomic binary predicate $\kappa XY$ is to be read as ’all $X$ is $Y$’.
From a formal point of view this is exactly what Aristotle is expressing and what Aristotle makes use of in the structure of the proofs in the Analytics. Aristotle certainly seems to assume the following reductive axiom (or rather definition) in the proofs:
$\forall A,B(\kappa AB\leftrightarrow\forall x(Ax\rightarrow Bx))$
Thus $\kappa$ is extensional. But Axiom (A) has to be tweaked as it is unlikely that Aristotle would have accepted $\kappa AA$.
We introduce an extensional second-order equality predicate
$A=_{e}B\leftrightarrow\forall xAx\leftrightarrow Bx$
and then replace $(A)$ with
$\forall A,B(\kappa AB\leftrightarrow\forall x(Ax\rightarrow Bx)\&A\neq_{e}B)$
It is well known, Aristotle seems to assume that all terms have non-empty extension (even elements of the such extensions are classed as ’non-being’). Thanks to ASOL we can state this directly as an axiom
$\forall A\exists xAx$
Another fundamental second-order binary predicate is $\mu XY$ which expresses ’some $X$ is $Y$’. We have the following axiom or definition
$\forall A,B(\mu AB\leftrightarrow\exists x(Ax\&Bx))$
which states that $\mu$, like $\kappa$, is extensional. Again we mention the problem that we have to exclude $\mu AA$ and this leads us to
$\forall A,B(\mu AB\leftrightarrow\exists x(Ax\&Bx)\&A\neq_{e}B)$
We can follow the approach of [14] and introduce a second-order monadic function to represent term-negation (de re negation) which we denote by $A^{c}$. But thanks to our ASOL framework we can now define it directly
$A^{c}=\lambda x\sim Ax$
Note that from (NEmp) we deduce that $\forall A\exists x\sim Ax$. Finally we can define $\hat{\kappa}XY\leftrightarrow\kappa XY^{c}$ and $\hat{\mu}XY\leftrightarrow\mu XY^{c}$. Alternatively we could introduce $\hat{\kappa}$ and $\hat{\mu}$ directly with the expected axioms, for instance
$\forall A,B(\hat{\mu}AB\leftrightarrow\exists x(Ax\&\sim Bx))$
We now present some evidence that the ASOL approach captures the structure of the metalogical proofs in the Prior Analytics.
Aristotle in 25a announces the conversion rule
$\forall A,B(\hat{\kappa}AB\rightarrow\hat{\kappa}BA)$
Let us pay close attention to how he proves it in the text. Aristotle’s argument is closely paraphrased as follows:
Let $A$ and $B$ be terms and assume that $\hat{\kappa}AB$. Assume per absurdum that $\sim\hat{\kappa}BA$. Then by definition of $\hat{\kappa}$, $\sim\forall x(Bx\rightarrow\sim Ax))$, that is
$\exists x\sim(Bx\rightarrow\sim Ax)$
were we note the implied knowledge of propositional logic. Using again this knowledge the above is considered equivalent to
$\exists x(Bx\&Ax)$
Now we argue that Aristotle is actually deploying the $\exists^{1}$ rule, so he considers a witness $C$ (we use $y$) and introduces the hypothesis $By\&Ay$. But by hypothesis $\hat{\kappa}AB$, that is $\forall x(Ax\rightarrow\sim Bx)$. But then using $\forall^{1}$ with $y$ we get
$Ay\rightarrow\sim By$
[…]
Aristotle’s goal in chapters 4 - 7 of the Prior Analytics is to find all valid sentences of the form (Syll). While the logically valid syllogisms can be seen as being deduced in second-order logic from the axioms (mainly (Econ)), invalid sentences are eliminated by exhibiting counterexamples. Metalogically we must be able to express that a certain sentence is not a syllogism, for example in 26a:
$\sim\forall A,B,C(\kappa AB\&\hat{\kappa}CB\rightarrow\kappa CA)$
which Aristotle takes to be equivalent to
$\exists A,B,C\sim(\kappa AB\&\hat{\kappa}CB\rightarrow\kappa CA)$
that is
$\exists A,B,C(\kappa AB\&\hat{\kappa}CB\&\sim\kappa CA)$
This in turn is shown by producing concrete witnesses. In this case $A=animal,B=man,C=stone$. Aristotle assumes as valid
$\kappa\,man\,animal\&\hat{\kappa}\,stone\,man\&\sim\kappa\,stone\,animal$
and his reasoning can be interpreted as applying repeatedly second-order $\exists^{2}I$ to obtain (b) which is then seen to be logically equivalent to (a). What we have seen exhibits non-trivial reasoning with quantifiers over non-monadic (second-order) predicate. A nice aspect of the ASOL intepretation is that we do not have to leave ASOL to formalize Aristotle’s counterexamples.
Thus second-order logic with natural deductive rules appear to be a suitable framework to interpret the metalogic, the argumentation in the Prior Analytics themselves.
There are other passages where Aristotle can be seen as using a mixture of universal and existential quantification over monadic predicate variables. In the Posterior Analytics Book I, chapter 15, Aristotle gives the definition of immediate predication which can be expressed directly in ASOL as
$\forall A,B(\hat{\kappa}^{\ast}AB\leftrightarrow\kappa AB\&\sim\exists C(\kappa AC\&\kappa CB))$
We now consider briefly how the modal syllogistic could be treated in the framework of ASOL. To this end we consider an additional second-order monadic function $\square X$. This takes a monadic predicate term $X$ and yields a monadic predicate term $\square X$. We posit the following axioms:
$\kappa\square AA$
$\kappa\square A\square\square A$
and the rule (K): from $\kappa AB$ we can derive $\kappa\square A\square B$. We also introduce the definition
$\sqcup A=\lambda x(\square Ax\vee\square A^{c}x)$
and call the resulting system ASOL. It seems very plausible that AML(see [14]) can be embedded into ASOL and all the corresponding theorems be derived.
In this section we have provided evidence that ASOL is a very natural candidate to express the content of the Analytics and its metalogical inference and use of counterexamples and have given examples of reasonings which can be seen very naturally as deploying both first and second-order versions of the introduction and elimination rules for existential and universal quantifiers as well as definitions and propositions in which these two types of quantifiers occur in a non-trivial way over (second-order) binary predicates.
Natural Deduction in Aristotle and Galen
In this section we focus specifically and in more detail on evidence for Aristotle’s conscious understanding and deployment of the first-order natural deduction quantifier rules and supplement our argument with the valuable evidence provided by Galen. From what we said in the previous section Aristotle’s deployment of the second-order quantifier rules $\forall^{2}$ and $\forall^{2}$ is evident and the same goes for $\exists^{2}$, for instance in his frequent use of examples with concrete terms. Our discussion in the next two sections about the Topics provides further evidence. The use of $\exists^{2}$ still needs to be investigated.
We argue that in Aristotle (and Galen) there is evidence for i) the conscious (philosophically reflected) use of the first-order introduction and elimination rules for $\forall$ and ii) that these rules were applied to multiple universal quantification over non-necessarily monadic predicates.
[…]
In the Organon and other works Aristotle presents us with evidence of complex and accurate propositional reasoning. For instance in how constructive and refutation topics are inferred from each other, as will be illustrated in the following sections. This fact cannot be explained by mere natural language competency, mathematical training or a spontaneous manifestation of Aristotle’s specific intellectual gifts. Rather there must have been a significant continuity between Aristotle’s or his disciple Theophrastus’ (who wrote more logical treatises than his teacher) and Chrysippus’ patently sophisticated formal system of propositional logic[4].
In the Analytics universally quantified sentences are considered which include non-monadic predicates. Such was the typical form of an axiom or principle. Consider the discussion of the axioms that need to be applied in the geometric proof in the Prior Analytics 41b13–22. One axiom mentioned can be written directly in ASOL as
$\forall xyzw(x=y\enspace\&\enspace z=w\rightarrow x-z=y-w)$
[…]
$\forall abc(Double\,ab\enspace\&Double\,bc\rightarrow Quadruple\,ac)$
[…]
We now argue i) and ii) for the case of $\forall$-introduction. Consider the remarkable passage from chapter 3 of Book II of the Topics (tr. W.A. Pickard):
Whereas in establishing a statement we ought to secure a preliminary admission that if it belongs in any case whatever, it belongs universally, supposing this claim to be a plausible one. For it is not enough to discuss a single instance in order to show that an attribute belongs universally; e.g. to argue that if the soul of man be immortal, then every soul is immortal, so that a previous admission must be secured that if any soul whatever be immortal, then every soul is immortal. This is not to be done in every case, but only whenever we are not easily able to quote any single argument applying to all cases in common euporômen koinon epi pantôn hena logon eipein , as (e.g.) the geometrician can argue that the triangle has its angles equal to two right angles.
To explain this passage we make use of the ASOL interpretation of the Topics developed in the next section. What Aristotle is first saying (in the intensional reading) is that in general we do not have ’induction’ from the species to the genus. If (essential) accident (or difference) $C$ applies to species $A$ of $B$ it does not follow that accident $C$ applies to $B$. That is, from $\gamma^{\star}AC\&\sigma AC$ it does not follow that $\sigma BC$. Thus if the human soul is immortal it does not follow that the soul is immortal (extensionally: every soul is immortal). Aristotle says that we have to assume not only $\sigma AC$ (and $\gamma^{\star}AB$) but also that $\gamma^{\star}AC\&\sigma AC\rightarrow\sigma BC$. Aristotle contrasts this situation to another situation in which euporômen koinon epi pantôn hena logon eipein - in which we are able to deduce (in a uniform way) a universal conclusion through a single argument, as the geometrician reasons about property of a triangle. We interpret this to mean, as seems natural, that this ’single argument’ is one involving a variable $X$ over which the universal conclusion is quantified. Thus the geometrician may start with ’Let $X$ be a triangle’ and arrive at the conclusion $X$ has its angles equal to two right angles and thereby view this as a proof that all triangles have their angles equal to two right angles. Thus we are lead to interpret Aristotle’s koinon epi pantôn hena logon eipein as a statement of the rule $\forall^{1}$ (and $\forall^{2}$).
It is also interesting to compare this to Proclus’ commentary of the first book of Euclid [16], p.162 which may be also interpreted as expressing the rule $\forall^{1}$:
Furthermore, mathematicians are accustomed to draw what is in a way a double conclusion. For when they have shown something to be true of the given figure, they infer that it is true in general, going from the particular to the universal conclusion. Because they do not make use of the particular qualities of the subjects but draw the angle or the straight line in order to place what is given before our eyes, they consider that what they infer about the given angle or straight line can be identically asserted for every similar case. They pass therefore to the universal conclusion in order that we may not suppose that the result is confined to the particular instance. This procedure is justified, since for the demonstration they use the objects set out in the diagram not as these particular figures, but as figures resembling others of the same sort (…)Suppose the given angle is a right angle. If I used its rightness for my demonstration, I should not be able to infer anything about the whole class of rectilinear angles; but if I make no use of its rightness and consider only its rectilinear character, the proposition will apply equally to all angles with rectilinear sides.
[…]
Everything that is in motion must be moved by something. For if it has not the source of its motion in itself it is evident that it is moved by something other than itself, for there must be something else that moves it. If on the other hand it has the source of its motion in itself, let AB be taken to represent that which is in motion essentially of itself and not in virtue of the fact that something belonging to it is in motion (…) a thing must be moved by something if the fact of something else having ceased from its motion causes it to be at rest. Thus, if this is accepted, everything that is in motion must be moved by something. For AB, which has been taken to represent that which is in motion, must be divisible since everything that is in motion is divisible. Let it be divided, then, at G. Now if GB is not in motion, then AB will not be in motion: for if it is, it is clear that AG would be in motion while BG is at rest, and thus AB cannot be in motion essentially and primarily. But ex hypothesi AB is in motion essentially and primarily. Therefore if GB is not in motion AB will be at rest. But we have agreed that that which is at rest if something else is not in motion must be moved by something. Consequently, everything that is in motion must be moved by something.
[…]
The previous observations lead us to conclude that ancient logic had at least the full expressive and inferential power of open theories. Open theories (sometimes called ”quantifier free” theories) are far from trivial as they include the theories of important algebraic structures such as real closed fields. A considerable portion of arithmetic can be formalized in the quantifier-free fragment of primitive recursive arithmetic [21]. It is also interesting to note (in our ASOL context) that in second-order logic we can define both first and second-order existential quantifiers in terms of universal quantifiers alone [12], p.67.
Similarly as for universal quantification we will attempt to show that there was a conscious use of the introduction and elimination rules for first-order existential quantification and specifically for such quantifiers applied to non-monadic predicates.
Consider again the proof of the conversion of the universal negation in the Prior Analytics 25a14-18 discussed in section 2 (tr. Hardie and Gaye):
First then take a universal negative with the terms A and B. If no B is A, neither can any A be B. For if some A (say C) were B, it would not be true that no B is A; for C is a B. But if every B is A then some A is B. For if no A were B, then no B could be A. But we assumed that every B is A
Aristotle’s argument (expressed in a rather terse and laconic style) can be read as follows: If A is not said of any B then neither is B said of any A. For assume that B is said of some A (Aristotle takes this immediately to be equivalent to it not being the case that B is not said of any A). Let this some A be $c$. Then $c$ is a $A$ which is also a $B$. But this contradicts the initial hypothesis that A is not said of any B. Hence by reductio ad absurdum we conclude that it is not the case that B is said of some A, that is to say, B is not said of any A.
The proof (which makes use of $\bot_{c}$) can be interpreted directly as follows in linearized natural deduction, in which we have condensed some logical steps and assumed basic Aristotelian premises in the Prior Analytics:
1. $\forall x(Bx\rightarrow\sim Ax)$ Hyp
2. $\sim(\forall x(Ax\rightarrow\sim Bx))$ Hyp
3. $\exists xAx\&Bx$ 2
4. $Ac\&Bc$ Hyp
5. $\bot$ 1,4
6. $\bot$ $\exists^{1}E$ 3-5
7. $\forall x(Ax\rightarrow\sim Bx)$ $\bot_{c}$ 2, 6
8. $\forall x(Bx\rightarrow\sim Ax)\rightarrow\forall x(Ax\rightarrow\sim Bx)$ $\enspace\rightarrow I$ 1,7
In 6 hypothesis 4 is discharged. Besides $\bot_{c}$ (which is common and consciously employed in Aristotle) the key ingredient is $\exists^{1}E$. The evidence points to $c$ being used in Aristotle’s proof exactly as $c$ in the natural deduction rule $\exists^{1}E$.
In 28a24-27 of the Prior Analytics Aristotle gives two proofs a certain syllogism in the third figure, the second one which employs a process called tô(i) ekthesthai poiein, byekthesis:
If they are universal, whenever both P and R belong to S, it follows that P will necessarily belong to some R. For, since the affirmative statement is convertible, S will belong to some R: consequently since P belongs to all S, and S to some R, P must belong to some R: for a syllogism in the first figure is produced. It is possible to demonstrate this also per impossible and by exposition (ekthesis). For if both P and R belong to all S, should one of the Ss, e.g. N, be taken, both P and R will belong to this, and thus P will belong to some R.
Recall that for Aristotle particular predication has existential import so we assume $\exists xSx$ as an axiom (NEmp). Aristotle’s second proof of this third figure figure (by exposition or ekthesis) can be read in detail as follows. Assume that all S is P and all S is Q. We have that there exists an S. Let $n$ be such an $S$. Then by the hypotheses we conclude that $n$ is $P$ and also that $n$ is $Q$. Thus some $P$ is $Q$.
The proof can be read as follows in linearized natural deduction:
1. $\forall x(Sx\rightarrow Px)\enspace\&\enspace\forall x(Sx\rightarrow Qx)$ Hyp
2. $\exists xSx$ Ax
3. $Sn$ Hyp
4. $Sn\supset Pn\enspace\&\enspace Sn\rightarrow Qn$ $\forall^{1}E$ 1
[…]
6. $\exists xPx\&Qx$ $\exists^{1}I$ 5
7. $\exists xPx\&Qx$ $\exists^{1}E$ 2-6
8. $\forall x(Sx\rightarrow Px)\enspace\&\enspace\forall x(Sx\rightarrow Qx)\rightarrow\exists xPx\&Qx$ $\enspace\rightarrow I$, 1, 7.
In 7 we discharged hypothesis 3. Again we have similar evidence for the deployment of the rule $\exists^{1}$ but this time in conjunction with $\exists^{1}$.
We propose that for Aristotle proof by ekthesis involved (essentially) the use of $\exists^{1}E$ in conjunction with other rules, perhaps mainly $\exists^{1}$.
Proclus considers constructions as a special type of proof. Unlike in modern ’classical’ mathematics Euclid does to revel in non-constructive existence theorems. A modern formalization of the first proposition of Book I would take the form $\forall x\exists yTxy$. This becomes through Skolemization
$\exists f\forall xT(x,fy)$
which is shown constructively by effectively producing such a $f$. So what is explicitly proven in Euclid is $\forall xT(x,fy)$, that for every segment $\overline{AB}$ the construction $f(\overline{AB})$ produces an equilateral triangle. It would seem that Proclus offers evidence that the ancients anticipated modern intuitionistic or constructivism interpretations of the existential quantifier666see[8] for further discussion on the constructivist interpretation of the existential quantification in Euclid. In appendix 3 we sketch how one might go about formalizing Euclid in ASOL in a way which seems natural and which also seems to fit within an intuitionistic or constructivist framework. in situations of multiple generality.
Sentences of the form $\forall x\exists yRxy$ are found in Aristotle. The (modern) concept of function corresponds to the case $\forall x\exists^{1}yRxy$ and is found embedded in Aristotle’s Physics. It appears in context of proofs in expressions in the grammatical form ’the $X$ of $Y$’ (i.e. the time taken of a given motion) wherein it is understood that $\forall x\exists^{1}yRxy$ is assumed to hold. The proofs sometimes deploy $\exists^{1}$ on such expressions and this takes the form ’let $A$ be the $X$ of $Y$’ (for example: let $A$ be the time taken by body $X$ in motion $Y$). In the Physics there is an implied functional correspondence, for a given body, between all the continua of motion, space, time and magnitude. A division of one of these corresponds to a division of the others. Aristotle typically presents arguments involving a body $A$ moving in time $T$ together with the construction: ’let $S$ be the magnitude moved by $A$ in time $T$’, i.e. ’the $S$ of $A$ in $T$’ which supposes something like $\forall xt\exists^{1}yMov\,xyt$. The principle manner in which such suppositions are used corresponds to the deployment of $\exists^{1}$ together with other rules - which is our proposed interpretation of ekthesis. This technical term can be read as a sign that for Aristotle the deployment of the natural deduction rules for the existential quantifier was conscious and reflected.
There are not many detailed mathematical proofs in the surviving Aristotelian corpus but the geometrically grounded proofs in the Physics and De Caelo et Mundi (see [6] for a good discussion of the mathematics of Aristotle) leave little doubt as to the presence of sentences with embedded quantifiers. For instance in Book VII of the Physics777241a25 in the proof we discussed that everything that moves is moved by something, hapan to kinoumenon hupo tinos anagkê kinesthai, Aristotle makes use of this axiom:
$\exists x(Stops\,xt\rightarrow Stops\,yt)\rightarrow\exists zMoves\,zyt$
where $t$ represents a given time:
In the second place that which is in motion without being moved by anything does not necessarily cease from its motion because something else is at rest, but a thing must be moved by something if the fact of something else having ceased from its motion causes it to be at rest.
As we will discuss in the next section, Aristotle’s theory of definition in the Topics shows traces of a more complex theory involving a theory of relations and in particular there must have been a theory of how definitions of relations were carried out. How would basic relations such as that of $x$ being the paternal grandfather of $y$ ($PGxy$) have been defined ? There is little doubt that Aristotle would have accepted the definition:
A being the paternal grandfather of B is A being the father of C and C the father of B for some C
This can be written
$PGF\,xy\leftrightarrow\exists z(Father\,xz\enspace\&\enspace Father\,zy)$
Even the most basic reasoning involving such relations would require something like Aristotle’s ekthesis and $\forall^{1}$. This we present as evidence that ekthesis was a fundamental tool of Aristotle’s logic. Let us give an example of one such argument and how Aristotle might have deployed ekthesis in a way strictly analogous to the examples from the Analytics discussed previously. Let us assume the axiom every man has a father $\forall y\exists xFather\,xy$. How would Aristotle have proven that every man has a paternal grandfather, $\forall y\exists xPGF\,xy$ It seems reasonable to suppose that this would have been done by rules for $\forall^{1}$ combined with ekthesis (according to our interpretation) in analogy to the proof 28a24-27 we analysed: by (in this case a repeated use of) a combination of $\exists^{1}$ and $\exists^{1}$. Given a man $m$, we instantiate the axiom to $\exists xFather\,xm$ and then assume via ekthesis that $Fcm$ and then repeat the process for $c$ instead of $m$ to assume $Fc^{\prime}c$ again via ekthesis. We then obtain $Fc^{\prime}c\,\&\,Fcm$ and finish the ektheses by two applications of $\exists^{1}$ to obtain $\exists zPGFzm$. By analogy from the Analytics Aristotle might have expressed this succinctly as follows:
Let $M$ be a man and let $c$ be $M$’s father. Then $c$ has a father and let $c^{\prime}$ be $c$’s father. Then there is some man having $c^{\prime}$ as father and who is the father of $M$, i.e., $c^{\prime}$ is the paternal grandfather of $M$. Thus $M$ has a paternal grandfather.
We conclude this section by presenting some evidence for the use of quantifier rules over non-monadic predicates and situations embedded quantifiers in Sextus Empiricus. Note that we are evidently not implying that the mere presence of expressions involving multiple generality implies that the ancients had a mechanism for dealing with inferences over such kinds of expression. Rather we are presenting evidence from Sextus of inferences involving multiple generality presented explicitly as inferences888such inferences, if correct, are no small achievement if one bears in mind how often even the most basic fallacies in propositional logic are committed by people who we would not expect to exhibit faulty reasoning. Arguably natural language competency does necessarily imply logical competency. Showing that multiple generality could have been expressed among the Stoics by a systematic grammatical regimentation as argued in [5] is certainly an important step towards arguing for the historical presence of a Stoic quantifier logic. . It is likely that in late antiquity and the Hellenistic era logic become more syncretic and that there was substantial borrowing between the different schools. Thus even if the passage of Sextus in question seems to primarily concern Stoic logic it is probable that it also reflects indirectly later Peripatetic logic as well in particular with regards to the hypothetical syllogism expounded by Galen. Consider Sextus’ Outlines II, v, 23 (discussing a view of Democritus)[19]:
Democritus says that a human being is what we all know. But as far as this goes, we shall not be acquainted with humans; for we also know dogs — and for that reason dogs will be humans. And there are some humans we do not know — so they will not be humans. Or rather, as far as this conception goes, no-one will be a human; for if Democritus says that a human must be known by everyone, and no human is known by all humans, then no-one will be a human according to him.
We read the deduction of the last sentence as follows: From the premises
$\forall x(Human\,x\leftrightarrow\forall y(Human\,y\rightarrow Know\,yx))$
$\sim\exists x(Human\,x\&\forall y(Human\,y\rightarrow Know\,yx))$
is concluded
$\sim\exists xHuman\,x$
This is a case of multiple generality involving non-monadic predicates. Presumably Sextus is deploying Stoic but maybe also Peripatetic rules to derive the conclusion. If we assume an Aristotelic argumentation then according to our interpretation it might have been carried out as follows. We assume that $\exists xHuman\,x$ then we use a strategy involving $\exists^{1}$. So let $a$ be an existing human,$Human\,a$. We instantiate (1) and through propositional reasoning obtain that
$Human\,a\&\forall y(Human\,y\rightarrow Know(y,a))$
we then apply $\exists^{1}$ to obtain a contradiction $\bot$ from (2). We conclude by applying $\exists^{1}$ to eliminate the hypothesis $Human\,a$ and conclude with $\bot_{c}$. This procedure would seem again quite analogous to the proof by ekthesis in Aristotle we examined.
There also an interesting argument in II, xii, 141:
If some god has told you that this man will be rich, then this man will be rich. But this god (I point, let us suppose, to Zeus) has told you that this man will be rich. Therefore, this man will be rich.[19]
. We interpret this as:
$\forall y((\exists xGodx\&ToldRich\,xy)\rightarrow BeRich\,y)$
$God\,a\&ToldRich\,ab$
therefore
$BeRich\,b$
The natural deduction proof of this inference involves essential use of both $\forall^{1}E$ and $\exists^{1}I$. Here we leave aside the possibility of a more complicated ’intensional’ interpretation of this argument as in [5], that is, a reading of the form ’a god $x$ told $y$ that $[\phi]$’. However [5] does present evidence that the Stoics did indeed consider such complex kinds of lekta.
The Logic of the Topics
Aristotle’s Topics is generally considered to be an early and indeed composite work, reflecting perhaps rather than specifically Aristotelian ideas, common dialectical procedures employed at the Platonic Academy. Indeed the intended purpose of the work is precisely to serve as a debating manual. It is also the part of the Organon that has received less attention from scholars. We hold that such a neglect is undeserved and that the Topics is quite a rich work and of considerable interest in its own right. Indeed it is impossible to do justice to all its aspects in the scope of the present paper. The Topics transcends the strictly ’logical’ concerns of the Analytics as grammatical, semantic, ontological, pragmatic and epistemic aspects equally predominate. From the semantic angle the concern with ambiguity and synonymity is almost an obsession, a fact that is likely explained by preoccupation with facing the rhetorical devices of the sophists. In this section we only consider some aspects of the purely logical content of the Topics. The Topics is a debate manual, but for a very restricted and formalized kind of debate, concerned with set of basic relations among terms [^9] Unfortunately Aristotle’s view of grammar and his theory of terms and concepts is only documented in a few scattered passages (but see chapter 4 of book IV which deals with relative terms and their governed cases.). We know more about Stoic grammar and its grammar-based logic. A basic linguistic entity is the sentence or protasis. A sentence is made up of two linguistic expressions connected by a certain propositional relation. The two expressions are called subject and predicate depending on their position in the propositional relation. Only noun phrases can be a subject but adjectival and verbal constructions can be predicates. These can becomes subjects after a certain syntactic transformation. The original concept of ’verb’ appears to have been that of predicate. A ’term’ oros is what can be a subject. However there is also a sense in which a verb is term to which is added a temporal aspect. The Stoic concept of kategorema seems to fit more exactly the concept of predicate. A predicate can be considered in itself, as unsaturated by any concrete subject. In Plato’s Sophist there is discussion concerning ’sentences’, ’nouns’ and ’verbs’, sentences being combinations of nouns and verbs, and all sentences must have a ’subject’. Plato gives the example ’Theaetetus, with whom I am now speaking, is flying’. This illustrates how Plato’s concept of ’noun’ corresponds to our ’noun phrase’ and very likely to the Aristotelic term or ’oros’ . (which the scholastics called predicabilia), the most important being the species-genus relation, the property relation, the relation of difference (which combined with the species-genus relation yields the further relation of definition) and the relation of accident [^10] It is not clear if this exhausts all the forms of predication. Indeed Aristotle often mentions predication without any further qualification and we are left to wonder if it assumed that this predication must fall within one of the classes above. Also ’accident’ is understood in various different ways.. The second-order predicates $\kappa$ and $\mu$ of the Analytics are a simplification and abstraction of these relations. Following [15] we hold that a topic is a sentence, generally in universally quantified form, expressing properties of these fundamental relations which can be used either to construct or refute a thesis involving one of the predicabilia. We hold that second-order logic is required to express the topics as well as the logical equivalences and inferences implicitly assumed to hold amidst related groups of topics [^11] In 100a Aristotle gives the definition of ’syllogism’. The genus is ’logos’. The difference is that in this logos certain premises being posited something different (heteron) from them occurs (sumbainei) out of necessity. We might think that a syllogism thus corresponds to our modern notion of logical ’rule’ (as in natural deduction or sequent calculi). The problem is that we would then have to reject rules such as $\Gamma,A\vdash A$ for $A$ is not different from the premises. Our position is rather that ’syllogism’ in this context corresponds not to the modern notion of ’rule’ but to a universally quantified second-order sentence as explained further ahead. . In our framework of ASOL the predicabilia are again second-order binary predicates which we denote by $\gamma$ (species-genus), $\iota$ (property), $\delta$ (difference) and $\sigma$(accident). A very important aspect is that they are no longer extensional, contrary to the case for $\kappa$ and $\mu$, that is, our second-order relations cannot be reduced to expressions in which second-order predicates do not occur. And at the same time relations of extension play an important role in many topics: we need to express that the extension of one predicate is strictly contained in another or that they are equal - and this justifies the second-order logic approach because this task can be carried out with the help of the first-order component of the logic. For example if we have $\gamma AB$ then the extension of $A$ is strictly contained in that of $B$, that is, $\forall xAx\rightarrow Bx$ and $\exists xBx\&\sim Ax$, but this last condition is not sufficient (contrary to the case of $\kappa$) to guarantee that $\gamma AB$. Also noteworthy is that the Topics is not restricted to terms corresponding to first-order monadic predicates, there are ’relation’ terms as well, corresponding to binary or higher-order first-order predicate terms. And we can ask what for Aristotle the ’extension’ of such as relation term would be.
The Topics (which has traces of being a compilation of diverse, not always consistent, texts written at different times) consists of eight books. Book IV is dedicated to $\gamma$, book V to $\iota$ and book VI to definition (including a treatment of $\delta$). Important material relating to $\sigma$ is scattered throughout the first two books. Book IV is the most approachable from a formal point of view and seems to also provide a transition to the Analytics. Let us now recapitulate on our fundamental second-order relations and state their corresponding basic axioms relating to extensions. For convenience we will in general write the topics as open formulas where the monadic predicate variables are assumed to be universally quantified. The most common form of a topic is
$\forall A,B(\phi(A,B)\rightarrow\bigcirc AB)$
[…]
$\forall A,B(\bigcirc AB\rightarrow\phi(A,B))$
For the genus-species relation (a second-order binary predicate $\gamma$) we have:
$\gamma AB\rightarrow\forall x(Ax\rightarrow Bx)\&\exists y(By\&\sim Ay)\quad\text{($\Gamma$)}$
We have the difference relation $\delta AB$ read as ’$A$ is a difference of genus $B$121212 A careful reading of the text (specially Book VI) reveals that Aristotle uses ’difference’ in three different ways: the difference of a species (relative to a genus or a definition), coordinate differences (of the same genus) and differences that ’belong’ to a given genus. We consider the third to be the most fundamental and define the two previous notions of difference in terms of this last one.. For property [^13] The topics of property set the stage for the topics of definition which is the culmination of the Topics. The notions deployed for property are basically the same one as those deployed for definition but is a more straightforward way. There is a much greater laxness in what can count as a property as compared to definition. We might say that property represents of a more archaic form of definition before the requirement of the a regimentation into a (proximate) genus and difference was adopted. The basic idea is that the property of a term $A$ cannot involve $A$ itself (non-circularity), cannot employ less known terms and cannot contain redundancies. $\iota AB$ read ’$B$ is a property of $A$’ we have
[…]
None of these implications are equivalences [^14] There is no doubt that Aristotle conceives of co-extensive terms which are intentionally different (that is, they are essentially different terms, not only syntactically different.). Likewise the terms ’One’ and ’Being’ are described as being predicated of all other terms and yet intentionally distinct..
We could also venture the following axiom for accident [^15] Aristotle distinguishes between separable and inseparable (or essential) accident. But this seems to make sense primarily for a predication of an individual. $\sigma AB$ read ’$B$ is an accident of $A$’:
$\sigma AB\rightarrow\exists x,y(Ax\&Bx\&Ay\&\sim By)\quad\text{($\Sigma$)}$
However the version of accident $\sigma^{\prime}$ in the passage in 120b suggests we need a modal version of ASOL. An accident is simply a predicate which can both be and not be predicated of an individual which we could also read in ASOL as
$\sigma^{\prime}A\leftrightarrow\exists x(\lambda x(\lozenge A\&\lozenge A^{c}))x$
Consider axiom ($\Gamma$). This is an example of a topic. It is neither a topic of refutation nor directly a topic of construction - but an intermediate kind, a make sure that so-and-so topic. But from ($\Gamma$) Aristotle infers a powerful and convenient topic of refutation:
$\forall A,B(\exists x(Ax\&\sim Bx)\rightarrow\sim\gamma AB)$
which is topic 5 of Chapter 1 listed in appendix 1. By what we have seen we can assume that such an inference could be considered as obtained via Aristotle’s deployment of natural deduction rules for first and second-order quantifiers as well as propositional equivalences.
Although the topics are usually listed in an ad hoc fashion it is clear that they are organized into groups closely knit together by logical equivalence or implication. We hold that second-order logic is required both to formulate the topics and (in its natural deduction form) to explain how Aristotle might have established the inferences and equivalence therein. We might call the working out of these implicit inferences the ’metalogic’ of the Topics. Our objective here is not to undertake an exhaustive formal interpretation of all the logical topics in ASOL. Rather we will go through a selection of key topics mainly from book IV but also from books V and VI which provide evidence for the second-order nature of the metalogic. Let us start with book IV and $\gamma$. We find therein topics which state directly or are stated to be inferred from $\gamma$ having transitive, symmetric and anti-reflexive properties.
$\gamma AB\&\gamma BC\rightarrow\gamma AC$
$\sim\gamma AA$
$\gamma AB\rightarrow\sim\gamma BA$
as can be verified in appendix 1. The first topic is clearly an ancestor of the Barbara syllogism of the Analytics. Perhaps $\mu$ descends for instance from $\delta$ since the same difference can operate on distinct genera so that $\delta AB$, $A$ is a difference of genus $B$ means that ’some $B$ are $A$’ ( topic (1) further ahead has certainly some connection to the syllogism Darii). Then there is the important tree-like property:
$\gamma AB\&\gamma AC\rightarrow(\gamma BC\vee\gamma CB)$
which is the first topic in chapter 2 of book IV. Aristotle refers to $\gamma AB$ also as $A$ participating in $B$ or $A$ receiving the the definition of $B$. Aristotle mentions important axioms in the context of the derivation of refutation or establishing topics. This gives us insight into the ’metalogic’ of the topics. For instance in 122b Aristotle infers a topic using the tree-like property and the topic can itself be represented as inference in sequent form:
$\forall AB(\gamma AB\&\gamma AC\rightarrow(\gamma BC\vee\gamma CB)),\gamma DE,\gamma DF,\neg\gamma EF\vdash\gamma DE$
which of course makes use of the ’transitive’ property of $\gamma$.
There is the defined concept of proximate genus:
$\gamma^{\star}AB\leftrightarrow\gamma AB\&\sim\exists C(\gamma AC\&\gamma CB)$
This furnishes us with an example of nested heterogenous quantifiers over non-monadic (second-order) predicates.
Another example: every genus must have more than one proximate species (a species of which it is the proximate genus)
$\gamma^{\star}AB\rightarrow\exists C(\gamma^{\star}CB\&C\neq A)$
Here we have an appearance of equality for monadic predicate terms, which we should add to our arsenal of second-order predicates - for we cannot reduce it to first-order extensional equivalence ($=_{e}$ of section2). We certainly have
$A=B\rightarrow(\forall xAx\leftrightarrow Bx)$
but not (as we argued previously) the reciprocal implication. We could tentatively posit that we have a second-order version of Leibniz’s law (see the topics in 152a which suggests particular cases of this law). In particular in book VII 152a31-33 we read (tr. W. A. Pickard):
Again, look and see if, supposing the one to be the same as something, the other also is the same as it: for if they be not both the same as the same thing, clearly neither are they the same as one another.
Aristotle appears to be stating that since $\sim(A=C\,\&\,B=C)\rightarrow\sim(A=B)$ then a valid topic is
$(A=B)\,\&\,(A=C)\rightarrow(B=C)$
The principle above about the plurality of immediate species furnishes us with important example of nested heterogenous quantifiers over non-monadic (second-order) predicates. Note that it follows from the axioms that the proximate genus of a term is unique (modulo the equality relation).
Axioms ($\Gamma$) and (I) are stated as a topics. So too are often the contrapositives of some of the topics considered above .
Our second-order binary predicates are also exclusive of each other. For instance we have the topic $\sigma AB\rightarrow\sim\gamma AB$ but this can be derived from ($\Sigma$) and ($\Gamma$). We also have that $\sim\Sigma AA$. There is a topic which seems to state that if $\gamma AB$ then $B$ cannot be the difference of a genus of $A$:
$\gamma AB\rightarrow\sim\exists C(\delta BC\&\gamma AC)$
If $\gamma AB$ then there must be some difference $C$ of $B$, $\delta CB$ which ’is said’ of $A$. Aristotle has a kind of unspecific general kind of ’predication’ wherein it is not obvious whether it is supposed to fall under one of the predicabilia or not. We take this unspecified predication to be just extensional containment $A\subset B\leftrightarrow\forall x(Ax\rightarrow Bx)$. We might express this topic as
$\gamma AB\rightarrow\exists C(\delta CB\&A\subset C)$
From this topic we derive the refutation topic:
$\forall C(\delta CB\rightarrow\sim A\not\subset C)\rightarrow\sim\gamma AB$
which is listed as number 11 of Chapter 2 in appendix 1. We find in addition the following interesting topic:
$\gamma AB\rightarrow\forall C(\delta CA\rightarrow\delta CB)$
If bipedal is a difference of mammal it is also a difference of animal.
We will not analyze here the question of supreme genera and infima species but the order-theoretic properties involving $\gamma$ could be easily expressed in ASOL though other implied properties such as the finiteness of ascending chains of subordinate genera would have to make use of an extension of ASOL such as Zalta’s Object Logic[22].
We have not addressed to the ontological matter of the different categories which feature in the Topics. For our purposes we can just remark that for the order on monadic predicate terms induced by $\gamma$ there can be in general more than one maximal element. And we can ask if it is correct to interpret our first-order individuals as individual substances. There is also a problem that Aristotle would seem to distinguish different kinds of situation for which $Ax$, for which a predicate term $A$ holds of an individual substance $x$, for example, if $A$ is predicated according to essence or as an accident. This would suggest that we need an extra second-order binary predicate with second-order and first-order arguments $\gamma^{i}Ax$ (for which $\gamma^{i}Ax\rightarrow Ax)$ to make such ontological predication distinctions of individuals. A general approach in ASOL the theory of predication in the The Categories might be something as follows (we will not discuss here the question of whether some of the other categories besides relation are in fact also relations: knowledge is a state of a subject but also, it seems, a relation). We postulate nine different categories of predication $\mathcal{P}^{i}Tx$ where we put $\mathcal{P}^{1}Tx\leftrightarrow Tx$ as corresponding to ousia while the others correspond to the other Aristotelian categories. The tenth will correspond to relation and be $\mathcal{P}^{10}R^{2}xy$ or $\mathcal{P}^{10}R^{2}xT$. For consider ’knowledge’ and ’knowledge of colours’. Knowledge corresponds to the state of a subject but also to a relation, and this relation can regard not one or many primary substances but other categories: if a subject $a$ knows the colour ’blue’ then this is not the same as knowing all blue substances.
A central role in the Topics is played be a kind of ’semantic’ term negation operator (opposite operator), which can be seen as operating within the proximate species of a given genus. This must be a second-order function applied (primarily) to monadic predicate terms. Its fundamental properties are
$\gamma^{\star}AB\rightarrow\gamma^{\star}A^{\bullet}B^{\bullet}$
$A=B\rightarrow A^{\bullet}=B^{\bullet}$
and we also assume $A^{\bullet\bullet}=A$.
For instance the opposite of the color white is black. A proximate species of a genus may not have an opposite; this situation which we could naturally express as $A^{\bullet}=A$. Noteworthy topics are
$\gamma AB\rightarrow\gamma A^{\bullet}B^{\bullet}$
and also
$B=B^{\bullet}\&\gamma AB\rightarrow\gamma A^{\bullet}B$
as can be seen in appendix 1.
In Book V we encounter the topic
$\iota AB\rightarrow\iota A^{\bullet}B^{\bullet}$
In order to formalize the theory of definition we cannot do without a binary function on monadic predicate terms which corresponds grammatically to a modifier in a noun-phrase, for instance an adjective attached to a head noun. We study in more detail the grammatical aspect of definition terms at the end of this section. We denote the function by means of concatenation and impose the axiom which is implicitly assumed by Aristotle:
$\forall x((AB)x\leftrightarrow Ax\&Bx)$
Since we do not want semantically meaningless constructs like ”green numbers” it makes sense to define a further second-order monadic predicate allowing us to express non non-empty extension (cf. also (NEmp) in section 2):
$AB\downarrow\leftrightarrow\exists x(AB)x$
so that
$\delta BC\rightarrow BC\downarrow$
The fundamental property of concatenation is that it lets a difference carve out a species:
$\delta DA\rightarrow\gamma(DA)A$
We can define definition [^16] Aristotle also uses the term ’logos’ for definition (cf. 141b3). according to Book VI as
$hAD\leftrightarrow\exists B,C(D=BC\&\delta BC\&\gamma^{\star}AC)$
We deduce immediately from $(I=)$ that $hAD\rightarrow(\forall xAx\leftrightarrow Dx)$. We may now formulate a fundamental axiom of Aristotelian logic, that every concept (in this case monadic predicate) has a definition:
$\forall A\exists D\,hAD$
From this and the axiom for our concatenation operation we can deduce the that immediate species of a genus have disjoint extensions:
$\forall AB\,\gamma^{\ast}AG\&\gamma^{\ast}BG\rightarrow A=B\vee(\sim\exists x(Ax\&Bx))$
It is important to note that equality here is not to be read syntactically but semantically. Thus $D=BC$ is to be read as ’the concept $D$ has as parts concepts $B$ and $C$’, the paradigm being ’rational animal’. We hold to be a fundamental problem whether a term can be said to be equal to its definition (with this notion of equality) and whether Leibniz’s law holds for this notion of equality. It could be that two distinct kinds of equality are required. It is true however that definition though the substitution of a term by its definition features in several topics171717152a-152b is very relevant for term equality. Aristotle announces topics which express the transitivity of $=$ as well as topics recalling Leibniz’s law. In 153a Aristotle seems to say that there is equality between a term and its definition but this is in itself not enough to show that it is a definition. The definiens must satisfy further conditions. The passage 140b32-141a1 is also interesting. Aristotle states that a definition is identical to the term it defines: esti de tauton tô(i) anthrôpô(i) zô(i)on pezon dipoun. However in 143a15-28 were Aristotle appears to admit that either giving the immediate genus and the difference or a non-immediate genus and all corresponding differences (the ’intension’) as both valid forms of definition which leads us to inquire exactly what the nature of the purported ’uniqueness’ of the definition. Does this means that both these forms are to be regarded as the same definition ?. From this last definition we can evidently derive (in ASOL) numerous refutation topics - and these are indeed topics in book VI. As Aristotle says: it is much easier to refute a definition than to establish one.
A standard view would be that if $hA(DC)$ then $C$ must be uniquely determined (this in fact follows from the above axioms). There is however a problem suggested by the passage 148b dealing with ’complex terms’181818this notion of ’complex term’ appears to be primarily grammatical and not logical or semantic as is our concatenation operation, however the precise relation between the two remains illusive as does the interpretation of the passage in question. Are we to understand the crucial non-circularity property of the definiens in semantic or syntactic terms ? But semantically does it even make sense to have a situation in which $A=AB$
Up to now ’difference’ was defined relative to a genus, that which carves out the species of a genus. We now wish to define what it means for a difference to be the difference of its species:
$\Delta AD\leftrightarrow\exists C(\gamma^{\star}AC\&\delta DC\&hA(DC))$
It is also interesting how definitions relates to the opposite operator. If we have a definition $DG$ of $A$ what can we say in general about the definition of $A^{\bullet}$ In 153b Aristotle says that there are three cases to consider for the definition for $A^{\bullet}$ depending on whether $G^{\bullet}=G$ or $D^{\bullet}=D$. The possibilities are either $D^{\bullet}G^{\bullet}$, $D^{\bullet}G$ or $DG^{\bullet}$. For example justice is virtue of the soul and its opposite injustice is the vice of the soul, soul having no opposite.
We now discuss Aristotle’s relation terms concerning which section 4 of book IV of the Topics is of particular interest. A natural approach is to represent relation terms by binary predicate terms in ASOL. But then we need a corresponding version of $\gamma$ for such arguments, there is no way to deal uniformly with the predicabilia when predicate terms of different arities are involved. This is also enforced by the topic in book IV section 1 (120b) which requires that species and genus belong to the same category, although Aristotle admits later in section 4 that there might be exceptions (124b) as for the case of ’grammatical’ (monadic) as a species of the genus ’knowledge’(relation). In our case the application of the aforementioned topic (stated also explicitly in (124b)) is simply that there must be two species-genus relations, one $\gamma$ for monadic predicates and another $\gamma_{2}$ for binary predicate variables. In a formalization we will be required to postulate analogous axioms as for $\gamma$, for instance:
$\gamma_{2}RS\&\gamma_{2}ST\rightarrow\gamma_{2}RT$
where $R,S$ and $T$ are binary predicate variables. The logic of relation terms is quite complex in Aristotle and bound up with Greek syntax in a way which renders interpretation difficult. One thing that is clear is that for a relation term $R$ we have its ’reciprocal’ or ’inverse’ which we can interpreted by a monadic second-order function on binary predicate terms thus $R^{\ast}$. A challenge is to interpret the passage in (124b) about the double ($D$) being the double of the half. This suggests first of all that Aristotle is considering one-to-one relations
$\forall x,x^{\prime},y,y^{\prime}(Rxy\&Rxy^{\prime}\rightarrow x=y^{\prime})\&(Rxy\&Rx^{\prime}y\rightarrow x=x^{\prime})$
One way of interpreting ’the double is double the half’ would be
$\forall xDx(\iota yD^{\ast}xy)$
so we can read the topic in (124b) as
$\gamma_{2}RS\rightarrow\forall xSx(\iota yR^{\ast}xy)$
or in its refutation version
$\sim\forall xSx(\iota yR^{\ast}xy)\rightarrow\sim\gamma_{2}RS$
In the passages that follow Aristotle discusses the necessary agreement between the case governed for $\gamma_{2}$. For example if $\gamma_{2}RS$ then if $R$ takes the genitive then so does $S$. More interesting from a logical point of view is the passage mentioning the number of complements, this suggests that Aristotle was conscious of the necessity of the general agreement between the arity of the predicate terms in the species-genus relation, so that we should consider a whole sequence of such relations $\gamma_{1}=\gamma,\gamma_{2},...,\gamma_{i},....$ where $\gamma_{n}$ serves for $n$-ary predicate variables. Aristotle uses the example of the trinary relations ’donation’ and ’debt’. If we look at the early and middle Platonic dialogues we find that relation plays a much larger and sophisticated role than in does in Aristotle’s Organon. The logic of relations does not appear to be very worked at our and was perhaps consciously and deliberately neglected. Indeed there are some natural questions that could be asked (both now and in antiquity). For example: how do definitions work for relations ? What are the differences and how are they joined to the respective proximate genera ? The search for the definition of relatives clearly figures in the Platonic dialogues, for example in the Lysis the definition of friendship is sought after. A definition of a relation can necessarily involve sentences of arbitrary complexity involving other relations which do not fall under the general scheme for monadic predicate terms of book VI of the Topics (at least for its most simple familiar cases). Indeed traces of such unorthodox definitions are to be founded scattered throughout this work191919for example the definition of Justice as Courage and Temperance (150a-151b).. Even the most basic relations such as that of ’being a paternal grandfather’ do not seem to be organized neatly into a proximate genus $+$ difference format. In the Lysis the binary relations of friendship $\Phi$ and love $L$ are considered. The following hypothetical definition is put forward:
$\Phi xy\leftrightarrow Lxy\&Lyx$
The Lysis (see appendix 2) deploys relatively complex statements and reasoning about relations - and none of the proposed definitions seem to fit the genus $+$ difference template. We can wonder why Aristotle did not take this more into consideration in the Topics and Analytics. However there are certain interesting passages which shed light on Aristotle’s view of relations. Consider the passage: 145a14-15: tôn gar pros ti kai hai diaphorai pros ti. Aristotle claims that differences of the relation ’science’ such as ’practical’, ’theoretical’ and ’productive’ are themselves relations. It is not immediately clear how to express this in ASOL (but see the discussion at the end of this section). The discussion in 142a is interesting for it appears to concede that the rule of defining the less know by the better known may fail for relatives such as ’half’ and ’double’. Such relatives must be defined in terms of their correlatives, the double in terms of the half (to diplasion hêmiseos, 142a26).
In appendix 1 we list some of the most important ’logical’ topics in Book IV including (but not limited to) those discussed in this section. We now investigate the structure of the noun phrases used as definitions
Classical Greek grammars do not give us much in the way of an account of the structure of noun phrases. Thus we will use the concepts employed in modern linguistic treatments of English grammar such as [9]. A noun phrase is usually considered as having the structure: determiner + premodifier + noun + postmodifer and relative clauses are included as a type of postmodifier. Obviously the case of being a pre- or postmodifier is not the same in ancient (Attic) Greek and modern English but we do not consider the position of a modifier expression relative to the head of the noun phrase to be of importance for our purpose, only the fact that it is a modifier. Definitions in the Topics are noun phrases and as such we must investigate their structure. Questions we wish to investigate in a future article include: how are we to understand grammatically the genus+difference template endorsed by Aristotle ? Does the genus always correspond to the head noun in the structure of the noun-phrase ? Does Aristotle allow definitions that do not fit this regimentation ? How do definitions of qualities differ grammatically from that of secondary substances ? What is the grammatical structure of the definitions of relatives ? For a term of a given category what are the constraints on the category of the difference ? And finally: is ASOL adequate to capture the logical content of all the grammatical forms of definitions considered by Aristotle ? In order to examine some of these questions it would be useful to list all the definitions (both accepted and rejected) discussed by Aristotle in book VI of the Topics but we list here only those definitions appearing in chapters 1 -5. We use the edition of the Greek Text from [2]. We present the definition noun phrase and then in parenthesis the noun of which the phrase is proposed of as a definition [^20] Aristotle often presents both the term and its definition in the accusative case following the construction: X (acc.) Y(acc.) if (or is) defined. We have endeavored in most cases to revert the term and the definition back to the nominative case. (whether accepted or not by Aristotle: most of the examples are rejected but for reasons which do not involve their grammatical structure).
  1. 1.
    139b20: agôgê eis ousian (genesis)
  2. 2.
    139b20-21: summetria thermôn kai psukhôn (hugieia)
  3. 3.
    139b32-34 : ametaptôton (epistemê), tithênê (gê), sumphonia (sôphrosunê)
  4. 4.
    140a3-5: ophruoskios (ophthalmos), sêpsidakes (phalaggion), osteogenes (muelos)
  5. 5.
    140b1-5: arithmos autos hauton kinoun (psukhê)
  6. 6.
    140b7-8: hugron prôton apo trophês apepton (flegma)
  7. 7.
    140a35 zôon pezon dipoun tetrapêkhun to epistêmês dektikon (anthrôpos)
  8. 8.
    140b23-24: zôon pezon dipoun tetrapêkhun (anthrôpos)
  9. 9.
    140b27-28 oreksin hêdeos (epithumia)
  10. 10.
    141a7 horistikên kai theôrêtikên tôn ontôn (phronêsis)
  11. 11.
    141a10 sterêsis tou kata phusin thermou (katapsuksis)
  12. 12.
    141a16 elattôsis tôn sumpherontôn kai dikaiôn (epieikeia)
  13. 13.
    141a19-20 epistêmê tôn hugieinôn zô(i)ôn kai anthrôpôn (iatrikê)
  14. 14.
    141a19-20 eikona tôn phusei kalôn kai dikaiôn (nomos)
  15. 15.
    141b21 epipedou peras (grammês), stereou peras (epipedos)
  16. 16.
    142b1 astron hêmerophanes (hêlios)
  17. 17.
    142b3-4 hêliou phora huper gês (hêmera)
  18. 18.
    142b8 to monadi meizon artiou (peritton)
  19. 19.
    142b11-12 arithmos ton dikha diairoumenon (artion), heksin aretês (to agathos)
  20. 20.
    142b24-26 to ekhon treis diastaseis (sôma), to epistamenon arithmein (anthrôpos)
  21. 21.
    142b30-31 to epistêmên tou grapsai to hupagoreuthen kai to anagnônai (grammatikê)
  22. 22.
    143a3 tou noson kai hugieian poêsai (iatrikê)
  23. 23.
    143a16-17 heksin isotêtos poêtikên ê dianemêtikên tou isou (dikaiosunê)
In (1) and (2) the genera seem to be agôgê and summetria while eis ousian and thermôn kai psukhôn would correspond to prepositional phrase postmodifiers in English. In (6) it appears that the head or genus is hugron which is affected by two modifiers, the adjective apepton and the non-finite clause equivalent to a relative clause prôton apo trophês were the infinitive (probably ’generating’) is omitted. While a common noun + adjective definition for secondary substances can be construed as a conjunction of monadic predicate constants $G$ and $D$ yielding a definition term $\lambda x(Gx\&Dx)$ the situation is more delicate for other kinds of postmodifiers. One approach is to consider the case in which the genus or head is relation. For instance agôgê (coming to be, arising) is modified to ’coming to be according to essence’. This perhaps would be $\lambda Ax\top$ where $\top$ is the supreme genera for ousia.
We cannot avoid mereology in (2): given a substance $a$ we need to form the sub-substances of the cold things in it $c$ and the warm things in it $w$. Could this be expressed by a term-forming operator syntactically analogous to set comprehension: $c=\{x\epsilon a:Cx\}$ and $w=\{x\epsilon a:Wx\}$ where we also have the analogue of $\cup$ ? If summetria is a relation $S$ then we interpret $Sxyz$ as meaning that $x$ has symmetry of $y$ and $z$. Thus our definition of hugieia (health) would look like
$H=\lambda y(Sy\{x\epsilon a:Cy\}\{x\epsilon y:Wx\})$
In (20) ’that which has three dimensions’ we have an example of a noun-phrase having as head a demonstrative pronoun with a relative clause postmodifier. The passage 140b32-141a1 is interesting. Aristotle’s point here seems to be that postmodifiers are not predicative so that a term may be predicated of a noun phrase (kata tou sumpantos) with a modifier employing the same term without redundancy (hapaks hê kategoria ginetai), for example in :a biped animal is biped. Finally we whould not fail to mention that ASOL is adequate to treat propositional attitudes through the use of second-order predicates involving 00-ary predicate terms (which correspond to propositions) and the $lambda$-construction for an empty list of variables as standard in SOL. For example $a$ knows that $Ab$ could be expressed as $Ka(\lambda(Ab))$. It would be interesting to look for instances of propositional attitudes in the Organon.
Formalizing book V
In this section we focus on book V and discuss some of the challenges facing the ASOL approach and its versatility in solving them.
It is probable that the elaborate theory of definition in book VI grew initially from the theory of property (idion) . Some textual evidence can be adduced in favor of this (cf. 154b). There are passages in which Aristotle relaxes all the complex structural requirements on the definiens term, even simple terms are allowed to be definitions; on the other hand a passage in 132a indicates that property must have a genus + specifications structure just like definition. How are we to understand that ’terrestrial bipedial animal’ indicates the essence of man but that (132a) ’tame (hêmeron) animal’ or ’animal gifted with knowledge’ somehow does not ? It might be tempting to view book V of the Topics, dealing with property, to be both among the oldest portions of the said work and perhaps even as having constituted an independent treatise. It is true that definition, even difference as well as predication ’according to essence’, are mentioned (and contrasted with property) in book V, but not in a detailed way, rather the passages in question could easily be viewed as later interpolations or revisions. We propose that book V be studied on its own in a more-or-less self-contained way.
In the most basic sense, property is related to co-extensionality as in $(I)$. But this is not all. Originally property, idion, seems to have been subordinated to pragmatic, epistemic purposes. Property was a kind of ’distinguishing mark’ (perhaps not unlike linga in the Indian tradition) which served to make an object known. Thus it is natural to suppose that the property is somehow conceptually simpler than that which it is a property of. For example as in certain marks distinguishing different species of plants.
For Aristotle co-extensionality is a necessary but not sufficient condition for being a property (cf. also 153a in book VII: equality is not enough for there to be definition). Thus being a property is intensional. If $\iota AB$ we cannot substitute co-extensional terms for $A$ or $B$ salva veritate. Indeed Aristotle’s predicate of terms ’ being better known’ (see 129b-130a) is also intensional. We denote this situation by using a second-order predicate $A\prec B$ to mean that the term $B$ is better known than $A$. $\prec$ is a transitive. We write $A\sim B$ if $A$ and $B$ are simultaneously known (this can be defined as $A\nprec B\&B\nprec A$). We meet the following rule in 131a: $A\sim A^{\bullet}$.
There is a series of syntactic and semantic conditions imposed on the property term. This theory involves implicit syntactic-grammatical notions which unfortunately are never made explicit in the surviving texts. Here are, in a modernized language, some of the conditions. For $B$ to be a property of $A$ it is required that:
  1. 1.
    the term $A$ does not occur in $B$ (non-circularity) (131a)
  2. 2.
    the term $B$ is different from the term $A$ (non-identity) (135a)
  3. 3.
    the term $B$ is not a conjunction $\lambda x(B_{1}x\&B_{2}x)$ in which $B_{1}$ or $B_{2}$ are already properties of $A$ (irreducibility)
  4. 4.
    $B$ is not a universal predicate or does not make use of universal predicates (non-redundancy)(130b)
  5. 5.
    the term $B$ does not make use of terms lesser known than $A$ (epistemic priority)(129b)
  6. 6.
    the term $B$ does not have repeated subterms (including via anaphora) (130b)
Perhaps 3 can be replaced with
3’.the term $B$ does not strictly contain a term $B^{\prime}$ which is a property of $A$
As a corollary of 5 we obtain that: $B$ cannot contain $A^{\bullet}$ (131a). 1 is actually strengthened by Aristotle to the requirement that $B$ does not even contain species of $A$.
Common modern systems of formal logic do not usually possess a mechanism to talk about their own syntax (at least not directly). And the conditions above do not seem to be primarily semantic (as our operation $AB$) but syntactic. But there is nevertheless a certainly a semantic aspect involved and we could try to formalize the semantic aspect in ASOL by means of our operation $(AB)$. For instance we could write
$A<B\leftrightarrow\exists C(A=BC)$
and $A\leq B$ for $A<B\vee A=B$.
and formulate weaker versions of some of the above topics as
  1. 1.
    $\iota AB\rightarrow A\nless B$
  2. 2.
    $\iota AB\rightarrow A\neq B$
  3. 3.
    $\iota AB\rightarrow\sim\exists C(\iota AC\&C<B)$
  4. 4.
    $\iota AB\rightarrow\sim\exists C(C\leq B\&\forall xCx)$
  5. 5.
    $\iota AB\rightarrow A\prec B$
It is difficult to formalize ’universal predicate’ for what kind of predication (out of the predicabilia) is this referring to ? Aristotle does not accept that a universal predicate can be a genus thus this can’t be predication understood as $\gamma$. The interpretation we take here is a weak version 4 referring to universal predicates according to extension. As our semantic concatenation operation arguably satisfies $AA=A$ it does not seem possible to express 6.
None of these conditions would we rejected, for instance, by a modern mathematician who wishes to characterize (but not necessarily define) a class of objects. Even if the ’better known’ relation is questionable, characterizations involving theory disproportionately more elaborate and distant than those of the the original class of objects will not always be welcomed.
Does it make sense to speak of a property of an individual ? How does one extrapolate from predication of individuals to a property of their species ? Aristotle introduces the construction: ’$B$ is said of all individuals $a$ such that $Aa$ qua $A$’ (132b). How are we to interpret this ? Aristotle in 132b speaks of the ’onoma’ and the ’logos’. But we could give a simple interpretation of the refutation topic in terms of extensions (which is equivalent to (I))
$\exists x(Ax\&\sim Bx)\vee(\sim Ax\&Bx)\rightarrow\sim\iota AB$
Aristotle’s own example involves $A$ as ’man’ and $B$ as ’living being participating in knowledge’. The individual $x$ is ’God’. Aristotle states that since $\sim Ax$ and $Bx$ it cannot be the case that $\iota AB$.
The following passage in (133b) is quite difficult to interpret: (W.A. Pickard translation):
Next, for destructive purposes, see if the property of things that are the same in kind as the subject fails to be always the same in kind as the alleged property: for then neither will what is stated to be the property of the subject in question. Thus (e.g.) inasmuch as a man and a horse are the same in kind, and it is not always a property of a horse to stand by its own initiative, it could not be a property of a man to move by his own initiative; for to stand and to move by his own initiative are the same in kind, because they belong to each of them in so far as each is an ’animal’ .
The subtext here seems to be that if $C_{1}$ and $C_{2}$ are species of a genus $G$ and $C_{1}$ has attribute $A_{1}$ and $C_{2}$ attribute $A_{2}$ then $A_{1}$ and $A_{2}$ must correspond ’in kind’. There seems to be analogy involved here. But why is ’moving’ to man analogous to ’standing’ for horse ?
Aristotle’s theory of opposition is difficult to fathom from a modern perspective. Opposition is not classical or intuitionistic negation. Rather it corresponds, as we saw, a second-order function defined on monadic predicate terms. But in fact there are many distinct species of opposition. $\sim Ax$ does not necessarily imply that $A^{\bullet}x$. The following is the fundamental topic for property and contraries:
$\iota AB\leftrightarrow\iota A^{\bullet}B^{\bullet}$
In section 2 we denoted Aristotle’s ’modern’ notion of term-negation by $T^{c}$. The associated topics are not surprising (136a). For instance $\iota AB^{c}\rightarrow\sim\iota AB$ and $\iota AB\rightarrow\sim\iota AB^{c}$.
Chapter 8 of book V deals with degrees of terms as well as with what might be called probabilistic logic. The aspect of degree of a term seems very much to act like an operator on terms which in ASOL is a second-order monadic function. Thus we can denote ’more’,’most’, ’less’ and ’least’ $T$ by $T^{+},T^{\triangle},T^{-},T^{\bigtriangledown}$. If a term does not admit degree then we assume that $T^{+}=T$, etc. The relevant topics can be condensed into
$\iota AB\leftrightarrow\iota A^{+}B^{+}\leftrightarrow\iota A^{-}B^{-}\leftrightarrow\iota A^{\triangle}B^{\triangle}\leftrightarrow\iota A^{\bigtriangledown}B^{\bigtriangledown}$
It does not seem that degree has a directly comparative aspect. Rather the degree is taken as relative to some implicit absolute standard. But for probabilistic degree the situation is otherwise. Let us write $\mathbb{P}\iota AB>\mathbb{P}\iota A^{\prime}B^{\prime}$ for ’it is more likely for $B$ to be the property of $A$ then $B^{\prime}$ to be the property of $A^{\prime}$ which is in reality a quaternary second-order predicate in the variables $A,B,A^{\prime},B^{\prime}$. Then Aristotle’s rather strange supposition can be expressed directly as
$\mathbb{P}\iota AB>\mathbb{P}\iota A^{\prime}B^{\prime}\,\&\,\sim\iota AB\rightarrow\sim\iota A^{\prime}B^{\prime}$
In 138a we see variants on the topics corresponding to the cases in which $A=A^{\prime}$ and $B=B^{\prime}$. There is also an analogous (quaternary) relation of equal likelihood:
$\mathbb{P}\iota AB=\mathbb{P}\iota A^{\prime}B^{\prime}\,\&\,\sim\iota AB\leftrightarrow\sim\iota A^{\prime}B^{\prime}$
as well as particular cases in which $A=A^{\prime}$, etc.
These topics are strange because Aristotle does not state clearly the intrinsic relationship between $A$ and $A^{\prime}$ or $B$ and $B^{\prime}$. The following difficult passage seems relevant:
The rule based on things that are in a like relation differs from the rule based on attributes that belong in a like manner, because the former point is secured by analogy, not from reflection on the belonging of any attribute, while the latter is judged by a comparison based on the fact that an attribute belongs.
Book V ends with the curious rejection of property given by superlatives - surely not the same superlative as $T^{\triangle}$ but rather something like a modifier of a noun, something of the form ’the Xest Y’. Aristotle says basically that such a definite description will not denote uniquely for different states of affairs.
[…]
Now Aristotle’s furnishes us with an interesting discussion of superlatives applied to pluralities. For instance the ’Spartans’ and ’Peloponnesians’ are both said to be the bravest of the Greeks. What do such plural superlative judgments mean ? How are we to make sense of Aristotle’s argument that the Spartans being the bravest and the Peloponnesians being the bravest only means that (the extension of) one term is contained in the other ?
The superlative in questions seems to be distinct and more extensional. We view Spartans and Peloponnesians not as species of ’Greek’ but as arbitrary terms having extensions contained in that of ’Greek’. We need a different concept of superlative. We assume that there is a partial order $<$ defined on the extension of $G$, that is, on individuals $x$ such that $Gx$. We write as before $A\subset B$ for $\forall x(Ax\rightarrow Bx)$. Let us posit that $S=sup^{2}_{<}G$ for $S\subset G$ with non-empty extension iff
$\forall x,y(Gx\&\sim Sx\&Sy\rightarrow x<y)$
We can now justify (in ASOL) Aristotle’s argument about the Spartans and Peloponnesians.
Theorem 5.1
If $S=sup^{2}_{<}G$ and $T=sup^{2}_{<}G$ then $T\subset S$ or $S\subset T$.
Proof.
Suppose that it is not the case that $T\subset S$ or $S\subset T$. Then (since we are assuming that terms $T$ and $S$ have non-empty extension) there is an $a$ such that $Ta$ and such that $\sim Sa$ and a $b$ such that $Sb$ and $\sim Tb$. But then by definition we must have both $a>b$ and $a<b$ and so $a=b$, a contradiction.
We end this section by illustrating how a passage of chapter 13 of book VI (about ’this and that’) might be interpreted (of interest because it relates to mereology). Let us define $A+B=\lambda x(Ax\&Bx)$. Suppose opposition does not satisfy the analogue of De Morgan’s laws but rather $(A+B)^{\bullet}=A^{\bullet}+B^{\bullet}$. Indeed it makes sense to say that the opposite of temperate and brave is intemperate and cowardly. Now suppose that we have defined a comprehension operation on individuals which allows us to form the pair $\{x,y\}$ which is also an individual as far as ASOL is concerned. Then we postulate that
$A\{x,y\}\leftrightarrow Ax\vee Ay$
With these premises we can easily derive Aristotle’s paradoxical conclusion that if we define justice to be temperance and courage then considering the pair of men in his example212121one is temperate but cowardly and the other is brave but intemperate. this pair will be both just and unjust. The elucidation of this paradox would lead to a closer examination of the ’mereology’ present in the Topics. In chapter 5 of book V there is a discussion of how property relates to the property of the parts of the thing it is a property of. Aristotle considers the special case of things which have like parts such as the sea. In this case the topic seems to be, if we denote the parthood relation by $A\epsilon B$:
$\iota AC\&B\epsilon A\rightarrow\iota BC$
hence ’being the largest body of water’ cannot be the property of the sea. This topic reflects a distinct mereological outlook from (M1).
Conclusion
We have shown how ASOL is the most natural framework to interpret the ’metalogic’ of the Analytics and the logic implicit in the Topics. The structure of the sentence and inferences involved confirms the competency of ancient logicians with regards to quantifier logic and multiple generality. If second-order logic is the most natural framework to read the Organon as a whole then this sheds new light of the whole standard narrative of the development of Western logic suggesting a much greater formal and conceptual continuity between ancient logic and Frege’s system of the Begriffsschrift which is usually equated to second-order logic with a term-forming extension operator. If Frege’s formal system has structural similarities with Aristotelian logic as argued for in this paper, his philosophy of logic and language has been argued (by Bobzien and others ) to have been influenced by the Stoics.
We now discuss in light of the present work the system of argumentation for which the Topics was written. In the Platonic Dialogues (particularly the early222222in a future paper we show the surprising second-order formalizability of the earlier Socratic dialogues as contrasted to the later ones. and middle ones) as well as in Aristotle there is the fundamental supposition that philosophical knowledge can be gained by finding the right definition of terms (under the proviso that the term in question is acknowledged to exist, that is, have non-empty extension.). This concern is not abandoned in the Analytics. The Socratic attitude is that the first step to wisdom consists in testing one’s beliefs which can lead to the valuable cathartic knowledge of the error of one’s beliefs. Thus a major part of philosophical activity involved the testing (i.e. attempt at refutation) of proposed definitions of terms - something which Aristotle acknowledges to be easier than finding correct definitions [^23] A massive amount of arguments were compiled during the eras of the late Academia and first Stoic schools, as testified by accounts of the lives of Chrysippus and Carneades.. In a modern framework how are we to understand (broadly speaking) the dialectical game of the Topics ? In ASOL in which we assume we have a finite collection of constants, functions and predicate symbols (first and second order) which represent the disambiguated vocabulary of a common language. The non-logical axioms we assume are $T$, the topics (or rather a minimal generating set for them ) and $K$ those representing common agreed upon knowledge. The game begins with Alice’s proposal of a definition of a certain term $hAD$. Then Bob basically has to try to derive a contradiction $T+K+hAD\vdash\bot$. The way the game was played apparently involved Alice confirming every move made by Bob, as attested in the Socratic dialogues (where it is rare for the ’opponent’ not to agree with Socrates’ moves). The ’moves’ can be $K$ axioms themselves or inferences, either purely logical or based on a topic. A nice example of this dialectical game in practice is given by Plato’s Lysis (see the ASOL formalization of a key section in appendix 2) although the types of definition considered do not fall into Aristotle’s scheme for monadic predicates terms outlined in book VI of the Topics. As discussed at the end of the previous section, it would be interesting to study the early and middle Platonic dialogues from a second-order perspective as in the present work so as to provide additional evidence for logical competency in dealing with multiple generality among the ancients.
We can of course question if this dialectical game could lead to actual progress in philosophy or to certain or plausible knowledge about reality. In particular the exact scope of $K$ seems to be vague and uncertain; the assumption of the possession of a perfectly disambiguated language is itself implausible (although it should be said that the Topics is acutely aware of this problem). However we should not go to the other extreme of discarding this kind of philosophical practice. Rather what was required is a preliminary radical epistemic, semantic and logical analysis in order to set up a relevant and meaningful system - and such concerns are in no way foreign to the mature Aristotle: things better known to us are not necessarily better known in themselves242424as we saw in section 5, the extra-logic second-order relation of ’being better known’ plays an important role in the Topics: if $A$ is a property of $B$ then it has to be better known than $B$. In book I,1 Aristotle says that a syllogism is a ’proof’ when the premises are ’true and primitive’ or epistemically depending on ’true and primitive’ premises. The adjective ’alethê’ is also translated as ’evident’. Aristotle then clarifies that evident and primitive is what is known through itself and not through another’. All ’scientific principles’ should be self-evident. Other propositions are ’endoksa’, translated sometimes as ’probable’..
As proposed in section 4, in order to gain a greater understanding of the logical, semantic and grammatical content of the Topics we propose the compilation of a list of all the definitions (both accepted and rejected) mentioned in this work and to pay close attention to their grammar and syntax and how this relates to their semantic and logical structure. We also need to clarify in greater detail the concept of predication deployed throughout the work and how it related to the concrete forms of predication given by the predicabilia. Additional indirect arguments for the adequacy of ASOL might be provided by showing that this is also the most natural approach to formalizing Euclid’s Elements, a prototype of which was surely known by Aristotle. We include a short discussion on the first proposition of Euclid in appendix 3.
Appendix 1: some key topics of book IV
Here we list the ASOL interpretation of several important logical topics in book IV. To this end we make use of the list compiled in [13][pp.691-694]. We divide the list into the corresponding chapters of book IV. We introduce an additional second-order predicate $\mathcal{C}AB$ to indicate $A$ and $B$ belong to the same category. We take $\gamma AB$ to encompass also what Aristotle calls ’$A$ participating of $B$’ and ’$A$ receiving the definition of $B$’ (121a). We use extensional containment $A\subset B$ for unspecified ’predication’.
Chapter 1.
  1. 1.
    $Ax\&\sim Bx\rightarrow\sim\gamma AB$ or $\gamma CA\&\sim\gamma CB\rightarrow\sim\gamma AB$
  2. 2.
    $\sigma AB\rightarrow\sim\gamma AB$ and $\sigma^{\prime}AB\rightarrow\sim\gamma AB$
  3. 3.
    $\sim\mathcal{C}AB\rightarrow\sim\gamma AB$
  4. 4.
    $\gamma BA\rightarrow\sim\gamma AB$
  5. 5.
    $\exists x(Ax\&\sim Bx)\rightarrow\sim\gamma AB$ (same as 1)
  6. 6.
    $\sim\gamma^{\star}BA\&\forall C(\gamma^{*}CA\rightarrow\sim\gamma BC)\rightarrow\sim\gamma BA$
  7. 7.
    $B\subseteq A\rightarrow\sim\gamma AB$
  8. 8.
    $Ax\&Ay\&(\sim Bx\vee\sim By)\rightarrow\neg\gamma AB$
Chapter 2
  1. 1.
    $\gamma AC\&\sim(\gamma BC\vee\gamma CB)\rightarrow\sim\gamma AB$ and in 4: $\gamma^{i}Ax\&\gamma^{i}Bx\rightarrow(\gamma AB\vee\gamma BA)$
  2. 2.
    $\gamma BC\&\sim\gamma AC\rightarrow\sim AB$
  3. 3.
    $\gamma BC\&\gamma CA\rightarrow\sim AB$ and also Ch.1,4
  4. 4.
    $Ax\&\sim Bx\rightarrow\sim\gamma$ and $Ax\&\gamma BC\&\sim Cx\rightarrow\sim\gamma AB$
  5. 5.
    $hBD\&A\not\subset D\rightarrow\sim\gamma AB$ and $hBD\&\gamma CB\&C\not\subset D\rightarrow\sim\gamma AB$
  6. 6.
    $\delta DA\rightarrow\sim\gamma DA\&\sim\gamma AD$ and $\Delta AD\rightarrow\sim\gamma DA\&\sim\gamma AD$
  7. 7.
    $\gamma BA\rightarrow\sim\gamma AB$ (cf. Ch.1,4)
  8. 8.
    included in 6. (122b) $\Delta AB\rightarrow A\subseteq B$.
  9. 9.
    included in 6. (123a) $\delta AB\rightarrow A\subseteq B$ (?)
  10. 10.
    $\gamma AB\rightarrow\sim\Delta AB$
  11. 11.
    $\forall D(\delta BD\rightarrow A\not\subset D)\rightarrow\sim\gamma AB$
  12. 12.
    (not logical)
  13. 13.
    $\exists x(Ax\&((\lozenge B^{c})x\vee(\delta DB\&Dx\&(\lozenge D^{c})x)))\rightarrow\sim\gamma AB$
Chapter 3
  1. 1.
    $B\neq B^{\bullet}\&\gamma AB^{\bullet}\rightarrow\sim\gamma AB$
  2. 2.
    $\sim\gamma AA$
  3. 3.
    $\sim(\exists C,D(\gamma^{\star}CB\&\gamma^{\star}DB\&C\neq D))\rightarrow\sim\gamma AC$
  4. 4.
    (not logical)
  5. 5.
    $B=B^{\bullet}\&\gamma AB\rightarrow\gamma A^{\bullet}B$ and $B\neq B^{\bullet}\&\gamma AB\rightarrow A\neq A^{\bullet}$ and $B=B^{\bullet}\&\gamma A^{\bullet}B\rightarrow\gamma AB$ and $B\neq B^{\bullet}\&\gamma A^{\bullet}B^{\bullet}\rightarrow\gamma AB$
  6. 6.
    (we cannot find this in the text)
  7. 7.
    $\gamma AB\rightarrow\gamma B^{c}A^{c}$
Chapter 4
9. $\sim\forall xSx(\iota yR^{\ast}xy)\rightarrow\sim\mathfrak{g}RS$ (124b)252525if $R$ is one-to-one.
Appendix 2: Plato’s Lysis as an example of ancient Dialectics
We apply ASOL to formalize the dialectics in 212-222 of the Lysis illustrating the naturalness of this approach. As in the Topics there is a second-order function which yields the ’opposite’ of a relation $R$ which we denote by $R^{\bullet}$. For instance if $Lxy$ means ’$x$ loves ’$y$’ then $L^{\bullet}xy$ is to be interpreted as $x$ hates $y$. We suppose with Plato that $R^{\bullet\bullet}=R$ and that $R=S\rightarrow R^{\bullet}=S^{\bullet}$. Also that $Rxy\rightarrow\sim R^{\bullet}xy$. We assume
$(Sxy\leftrightarrow Txy)\leftrightarrow(S^{\bullet}xy\leftrightarrow T^{\bullet}xy)$
We use a Zalta-type approach to Platonic forms whereby we gather one or more predicates into an abstract first-order object. Thus for instance $[X:X=\Phi]$ would be the singleton $[\Phi]$ which represents the abstract object encoding only $\Phi$ (see [22] for more information about the concept of encoding considered here).
We fix a subset $\mathcal{C}$ of all sentences of ASOL which we call sentences of common assent together with a set $\mathcal{T}$ of sentences considered to be true.
The Lysis concerns the definition of the binary relation $\Phi xy$ ($y$ is $x$’s friend). We are given a relation $Lxy$ ($x$ loves $y$). We now attempt to formalize the essential structure of the dialogue in 212-222.
Socrates begins by inquiring concerning the logical relationship between $\Phi$ and $L$. Lysis assents to the first hypothesis for a definition
$1.\enspace Lxy\enspace\&\enspace Lyx\leftrightarrow\Phi xy\quad\text{Hyp}$
Socrates then makes the deduction
$2.\enspace\sim Lyx\rightarrow\sim\Phi xy\quad\text{1}$
and brings out several sentences of $\mathcal{C}$ such as
$3.\enspace\exists x,y\,Man\,x\enspace\&\enspace Horse\,y\enspace\&\sim Lyx\enspace\&\enspace Lxy\enspace\&\enspace\Phi xy$
and $\exists x.Man\,x\enspace\&\enspace\Phi x\,wisdom\enspace\&\sim L\,wisdom\,x$ which refute 1 and thus we arrive at the conclusion
[…]
Then Socrates puts forward another hypothesis:
$5.\enspace Lxy\leftrightarrow\Phi xy\quad\text{Hyp}$
Applying (Op) we get
$6.\enspace L^{\bullet}xy\leftrightarrow\Phi^{\bullet}xy$
But clearly $\exists x,yLxy\enspace\&\enspace L^{\bullet}yx$. Hence we deduce that
$7.\enspace\exists x,y\,\Phi xy\enspace\&\enspace\Phi^{\bullet}yx$
There will be people who are friends of their enemies and enemies of their friends. The negation of 7 is in $\mathcal{C}$ and so hypothesis 5 is also rejected and we have
[…]
Socrates then makes an appeal to the wisdom of the poets and brings forth the predicates $Hxy$(likeness) and $Kx$ (good) alongside their opposites. For $H$ opposition and negation seem to coincide.
Likeness can be extended to $\mathfrak{H}xyz$ which is to be interpreted as ’likeness of $x$ to $y$ according to $z$’ and thereby resolve contradictions. So too for other relations. The following it put forth
$9.\enspace Hxy\leftrightarrow\Phi xy\quad\text{Hyp}$
Then follows interesting propositions involving $H$:
$10.\enspace\sim Hxx\rightarrow\forall y\sim Hxy$
$11.\enspace K^{\bullet}x\rightarrow\sim Hxx\enspace\&\enspace Kx\rightarrow Hxx$
From these propositions (assumed to be part of $\mathcal{C}$ or $\mathcal{T}$) Socrates then deduces (assuming $Kx\enspace\&\enspace Ky\rightarrow Hxy$) that
$12.\enspace Kx\enspace\&\enspace Ky\rightarrow\Phi xy\enspace\&\enspace K^{\bullet}x\enspace\&\enspace K^{\bullet}y\rightarrow\sim\Phi xy$
So bad people cannot friends.
Let $Pxyz$ mean $x$ can do $z$ to $y$. Socrates assumes
$13.\enspace Hxy\rightarrow\forall z(Pxyz\rightarrow Pxxz)$
Let $\mathfrak{K}xy$ be defined as $\exists z\sim Pxxz\enspace\&\enspace Pyxz$. The axiom of the self-sufficiency of the good is:
$Kx\rightarrow\sim\exists y\mathfrak{K}xy$
Socrates accepts the pragmatic condition for friendship
$14.\enspace\Phi xy\rightarrow\mathfrak{K}xy$
from which we deduce
$15.\enspace\Phi xy\rightarrow\sim Hxy$
and hence hypothesis 9 is refuted
[…]
Next, inspired by Hesiod, Socrates proposes:
$17.\enspace\sim Hxy\leftrightarrow\Phi xy$
Socrates then draws the obvious conclusion from 17:
$18.\enspace\Phi[\Phi][\Phi^{\bullet}]\enspace\&\enspace\Phi[\Phi^{\bullet}][\Phi]$
Friendship is friends with enmity and likewise justice is friends with injustice and so forth. Since this must be rejected we get
[…]
There can be a middle ground or mixture between opposites or a least a ground which belongs to neither of the opposites. Socrates proposes things which are neither good nor evil $\sim Kx\enspace\&\sim K^{\bullet}x$.
We denote this predicate by $Mx$. There are thus three classes of beings
$20.\enspace\forall x(Kx\vee K^{\bullet}x\vee Mx)$
We can use these classes to propose a definition of the form
$\mathbb{P}x\enspace\&\enspace\mathbb{Q}x\rightarrow\Phi xy$
where $\mathbb{P}$ and $\mathbb{Q}$ are among $K,K^{\bullet}$ and $M$. Since $x$ such that $K^{\bullet}x$ is ruled out from entering into $\Phi$ it follows from 15 and (K) that the only possibility is
$21.\enspace Mx\enspace\&\enspace Ky\rightarrow\Phi xy\quad\text{Hyp}$
Let $X$ be a $n$-ary predicate variable and $Y$ a $n+1$-ary predicate variable. We introduce the second-order relations $X\prec_{n}Y$ which must satisfy the axiom
$X\prec_{n}Y\rightarrow Xx_{1}...x_{n}\leftrightarrow\exists yQx_{1}...x_{n}y$
We consider $\Phi\prec_{2}\Phi^{+}$ in which $\Phi^{+}xyz$ is interpreted as $y$ is friend to $x$ for the sake of $z$. Socrates states that
$22.\enspace\Phi^{+}xyz\rightarrow\Phi xz$
and that the primary instance of $\Phi xy$ we are interested in is that in which $\Phi^{+}xyy$.
Apparently Socrates rules out an endless sequence $\Phi^{+}xya_{1},\Phi^{+}x,a_{1},a_{2},\Phi^{+}xa_{2}a_{3},...$ as well as a cycle such as $\Phi^{+}xyz,\Phi^{+}xzy$. Let $\mathfrak{F}xy$ be by definition $\Phi^{+}xyx$.
We also define $\Phi\prec_{2}\Phi^{a}$ where $\Phi^{a}xyz$ expresses that $z$ is a cause for $\Phi xy$ in the way that sickness is the cause of a doctor being a friend to a man. Socrates states that
$23.\enspace\Phi^{a}xyz\rightarrow K^{\bullet}z$
But
$24.\enspace\exists x,y,z\sim K^{\bullet}z\enspace\&\enspace\Phi^{a}xyz$
so 23 must be rejected.
There is the predicate $Bxy$, $y$ belongs to $x$. Socrates proposes
$25.\enspace Bxy\rightarrow\Phi xy$
but
$26.\enspace Kx\rightarrow(Bxy\rightarrow Ky)$
$27.\enspace Mx\rightarrow(Bxy\rightarrow My)$
$28.\enspace K^{\bullet}x\rightarrow(Bxy\rightarrow K^{\bullet}y)$
Hence since there is a $y$ such that $Bxy$ and $K^{\bullet}x$ then by $24$ we will have $\Phi xy$ which contradicts 12. Hence we reject 25.
Appendix 3: the first proposition of Euclid
Here we propose how a minimal fragment of Euclid (enough to formulate and prove the first proposition of the first book) might be formalized in ASOL.
Individuals correspond to points in the plane. There are no magnitudes as primitive sorts. Our primitives are $\Lambda Lxy$ for $L$ a monadic predicate variable, $\Sigma CxL$ for $C$ and $L$ monadic predicate variables and $A<B$ for $A$ and $B$ monadic predicate variables. The intended meaning of $\Lambda$ is: $L$ is a line with extremities $x$ and $y$, of $\Sigma$: $C$ is a circle with center $x$ and of radius equal to the length of line $L$ and $A<B$: line $A$ is shorter than line $B$. We employ, alongside standard equality for individuals, an extensional second-order equality on monadic predicate terms defined as
$A=B\leftrightarrow\forall x(Ax\leftrightarrow Bx)$
We define $A\equiv B\leftrightarrow A\nless B\&B\nless A$. This is to be interpreted as $A$ and $B$ having the same length (when lines) and should not be confused with extensional equality.
We assume that we can form ’intersections’ and ’unions’ and ’complements’ of monadic predicate terms in the expected way using the $\lambda$. For instance
$A\cup B=\lambda x(Ax\vee Bx)$
We can now define a triangle:
$\Delta Fabc\leftrightarrow a\neq b\neq c\&(\forall L_{1}L_{2}L_{3}(\Lambda L_{1}ab\&\Lambda L_{2}bc\&\Lambda L_{3}ca\rightarrow F=L_{1}\cup L_{2}\cup L_{3})$
and equilateral triangle
$\Delta^{E}Fabc\leftrightarrow\Delta Fabc\&(\forall L_{1}L_{2}L_{3}(\Lambda L_{1}ab\&\Lambda L_{2}bc\&\Lambda L_{3}ca\rightarrow L_{1}\equiv L_{2}\equiv L_{3})$
We assume the basic axioms
$\forall xy\exists L\Lambda Lxy$
$\forall xR\exists C\Sigma CxR$
$\forall xRCy\,\Sigma CxR\&Cy\rightarrow(\forall L\,\Lambda Lyx\rightarrow L\equiv R)$
and the axiom which Euclid took apparently for granted:
$\forall xyC_{1}C_{2}R(\Sigma C_{1}xR\&\Sigma C_{2}yR\&\Lambda Rxy\rightarrow\exists zC_{1}z\&C_{2}z)$
We now can prove
Theorem 6.1
$\forall xyL\,\Lambda Lxy\rightarrow\exists zT\,\Delta^{E}Txzy$
The proof in ASOL is surely shorter and more natural, closer to Euclid’s own proof, than a more convoluted first-order approach.
Appendix 4: the rules of ASOL
Prawitz[12] gives two versions of second-order logic (SOL) and ASOL can be built upon either of these versions. Here we only consider version 2 and adapt the notation to the use of the present paper. We define formula, individual term and $n$-ary predicate term as usual in second-order logic by simultaneous recursion as in [12]. Individual variables (which are individual terms) are denoted by $x,y,z,..$ and $1$-ary (monadic) predicate term variables (which are $1$-ary predicate terms) are denoted by $A,B,C,...S,T,...$. We denote a $n$-ary predicate term variable by $P^{n},X^{n},...$. Given a formula $\phi$ we can also form the $n$-ary predicate term $\lambda x_{1}....x_{n}\phi$. In the rules ahead $t$ denotes a variable term and $\mathfrak{T}^{n},\mathfrak{P}^{n}$ a $n$-ary predicate term. The language of ASOL is obtained by adding to SOL second-order predicate symbols $\alpha,\beta,...$ and second-order functions $f,g,...$. Each symbol is equipped with a signature which specifies its number of arguments (at least one) and for each argument whether it is an individual argument or a predicate argument, in which case the arity is also specified. For example $\gamma XY$ of section 4 is specified by having two predicate arguments. Note that if the arguments are all individual then the second-order predicate can be identified with a predicate term constant.
The rules are as follows:
[…]
the four first-order quantifier rules:
[…]
where in $\forall^{1}I$ $a$ must not occur free in any assumption on which $\phi$ depends and in $\exists^{1}I$ $a$ must not occur free in $\psi$ or in $\exists x\phi$ or on any hypothesis on which $\psi$ depends except for $\phi^{x}_{a}$. Here $\phi^{t}_{s}$ denotes the result of substituting $t$ by $s$ in $\phi$ [^26] Here we need not follow Prawitz’s terminological distinction between variables and parameters which correspond to bound and free variables..
We have the analogous second-order quantifier rules
[…]
with analogous provisos.
Finally there are introduction and elimination rule for $\lambda$-predicate terms concerning which the reader may consult [12].